OpenStudy (anonymous):

Which of the following equations is equivalent to 1/4x - 1/2y = 8? 2x - 4y = 8 2x - 4y = 64 4x - 2y = 64 C??

OpenStudy (tkhunny):

None of those. Here's some that are: 2x - 2y = 16 4x - 4y = 32 8x - 8y = 64

OpenStudy (anonymous):

those werent my choices

OpenStudy (tkhunny):

Well, that was my first comment. You need more choices or a "None of the Above". The choices given are absolutely NOT equivalent.

OpenStudy (anonymous):

THOSE ARE NOT THE CHOICES THE ONES ARE GAVE ARE THE ONLY ONES

OpenStudy (anonymous):

WAIT!!! that looks different... it said x-y=8... now it has a fraction

OpenStudy (tkhunny):

What are you not understanding. None of the answers given is equivalent. There is nothing you or I can do about it. Sometimes questions are written incorrectly. Let's get over it and move on, shall we? A: No. B: No. C: No. Where does that leave us? It leaves us with the next problem. You could check to see if you copied the problem correctly. That could be the problem.

OpenStudy (anonymous):

how do you get to 64 from 8?

OpenStudy (anonymous):

as in 64/8-?

OpenStudy (anonymous):

LOOOOK AT THE PROBLEM AGAIN AND MAYBE MAYBE YOU GOT IT WRONG

OpenStudy (anonymous):

STOP SCREAMING if you want help you need to be considerate

OpenStudy (anonymous):

its -3

OpenStudy (anonymous):

IM NOT SCREAMING MY BUTTONS ARE STICKY

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

ok then get off caps lock since it wasnt like that in the beginning...

OpenStudy (anonymous):

I CANT IT WILL GO AWAY IN A WHILE THE BUTTONS ARE STICKY.....SORRY

OpenStudy (anonymous):

what you need to do is find out how many 8's go into 64 .... then multiply that number why the entire equztion... you will get it then

OpenStudy (tkhunny):

Ah, I see we now have a new problem statement, rather than the x - y = 8 where we started. Very good. The correct problem statement often provides better solutions that the wrong one. 1/4x - 1/2y = 8 So much better. Multiply by 4 x - 2y = 32 Nope. Multiply by 2 2x - 4y = 64 There it is!!!!!! Looks like "B"! Yea.

OpenStudy (anonymous):

but i will agree when i first read the i saw x-y=8 ... in which case non of them did apply...

OpenStudy (anonymous):

THANKSSSSSSS